A fox locates its prey under the snow by slight sounds rodents make. The fox then leaps straight into the air, and then burrows its nose into the snow to catch its meal. if a fox jumps up to a height of 83cm. calculate how long the fox is in the air. Ignore air resistance.
Answer #1

v2=u2+2as , v is final velocity, u is initial velocity, a is acceleration , s is displacement

here v=0, a=-9.8 m/s2 ,s=0.83 m

u=4 m/s

v=u+at

v=0, u=4 m/s , a=-9.8 m/s2

time=t=0.41 s